### v1

# Towards AI-Assisted Synthesis of Verified Dafny Methods

Abstract.
Large stochastic language models show great promise in many domains, including programming. A promise is easy to make but hard to keep, and language models often fail to keep their promises when applied to programming, generating erroneous code. One promising avenue to keep models honest is to have them generate code in a language that supports formal verification: if and when that is adopted, the model would provide proof along with the code, and that proof would be automatically verified. Unfortunately, existing large language models show a severe lack of proficiency in verified programming languages.
In this paper we demonstrate how to improve two pretrained models’ proficiency in the Dafny verified programming language. Using 178 programming problems from the MBPP dataset, we prompt two contemporary models (GPT-4  and PaLM-2 ) to generate methods in Dafny. We use three different types of prompts: a direct contextless prompt, a second one that includes a signature of the method and test cases, and a third one that decomposes the problem into steps and includes dynamically chosen similar examples. Our results show that GPT-4 is better than PaLM-2 , but that, in both models, the third prompt greatly improves the success of the generation task with respect to the direct prompt. With the third prompt, GPT-4 was able to generate verified (and human-evaluated) Dafny methods in 58% of the cases, while the first prompt generated verified (and human-evaluated) methods in only 19% of the cases. Surprisingly, the second prompt had the worst performance, with only 10%.
One tangible contribution of our work is a collection of 153 MBPP problems that are implemented and formally verified in Dafny, 50 of which were written by us and 103 were automatically synthesized by GPT-4 . Additionally, our results demonstrate that the benefits of formal program verification (proof of correctness) are now within reach of large stochastic language models used to generate code. These results also demonstrate that program verification systems can likewise benefit from incorporating large language models, whether to synthesize code wholesale, to generate specifications, or to construct internal verification annotations such as loop invariants, that are hard for programmers and verification tools to find directly. (e.g. legal arguments, transport signaling, structural engineering, etc.) where solutions must be correct, and where that correctness needs to be verifiable by existing formal tools, or explained to (and understood by) designers and end-users.
Program Synthesis, Dafny, LLM, Program Verification
copyright: acmlicenseddoi: 10.1145/3643763journalyear: 2024submissionid: fse24main-p653-pjournal: PACMSEjournalvolume: 1journalnumber: FSEarticle: 37publicationmonth: 7ccs: Software and its engineering Formal software verification
1. Introduction
Formal verification is increasingly being adopted to support the development of high-quality, provably correct software. High-stakes domains, such as security-sensitive systems, cryptographic libraries, aerospace systems, and embedded software in medical devices, look to formal verification for correctness guarantees.
For instance, CompCert C (CompCert, 2023) is a formally verified C compiler that won the ACM Software Systems award in 2021;  [^1] https://awards.acm.org/software-system in a comprehensive study (Yang et al., 2011), researchers found no bugs in the CompCert C compiler compared to GCC (GCC, 2023) and LLVM (LLVM, 2023) toolchains. This study motivated Airbus to adopt CompCert C to help ensure safety and enhance aircraft performance (França et al., 2011). The seL4 project (Klein et al., 2009), awarded the ACM Software System Award in 2022, resulted in a formally verified high-assurance, high-performance operating system microkernel employed to protect an autonomous helicopter against cyber-attacks (seL4 Project, 2023). In order to secure communication, both Chrome and Android use formally verified cryptographic code (Erbsen et al., 2020). Similarly, Mozilla incorporated its verified cryptographic library for Firefox performance improvement (Jacobs and 2020, 2023).
Formal code verification consists of two parts that go hand-in-hand: formal specification of software properties, and automated, or semi-automated, verification of those properties. Over the past 50 years, verification has had a couple of major breakthroughs, first with the development of interactive theorem provers in the 1960s (Nederpelt et al., 1994), and then, at the turn of the millennium, with the development of Satisfiability Modulo Theory (SMT) solvers (Barrett et al., 2021). Unfortunately writing program properties and proofs is still a creative, manual process that requires significant effort, experience, and expertise. Formal specification languages are closer to mathematics than to regular programming languages, and necessarily incorporate subtle concepts based on classical or constructive logics. For example, writing the proofs for the seL4 microkernel was an eleven person-year effort of a group of experts (Murray et al., 2013). Another example: the verification code of the CompCert C compiler is more than three times the size of the original compiler code itself; it took three person-years to write the complete verification code, which is approximately two lines of code in a day (Leroy, 2009). The fact that these systems have received important awards is a testament to how exceptional these efforts are.
In the universe of languages and systems for formal code verification, Dafny (Microsoft, 2023a) stands out as having made a significant effort towards usability by programmers, rather than mathematicians. Dafny is a strongly typed imperative programming language with functional and object-oriented features that supports code verification via Hoare Logic (Hoare, 1969) assertions, preconditions, postconditions, and invariants (aka design by contract). Although algorithmic code in Dafny is similar to many other programming languages, writing the formal specifications and auxiliary verification assertions is still difficult (Faria and Abreu, 2023; Noble et al., 2022).
Over the past year, stochastic Large Language Models (LLMs) have been revolutionizing both the software industry and research in software engineering. LLM-assisted tools such as GitHub-Copilot (Copilot, 2023) and Amazon CodeWhisper (CodeWhisperer, 2023) have been accelerating several tasks of software development such as code generation, transformation, summarization, documentation, code review, program repair and synthesis. Although their performance is impressive for straightforward tasks in the most popular programming languages, given more complex problems or less familiar languages (such as Dafny), LLMs tend towards producing ”hallucinations” [^2] Respectable software engineers would call hallucinations ”faults”, ”errors”, or ”bugs”. (McKenna et al., 2023; Galitsky, 2023; Salvagno and Taccone, 2023; Azaria and Mitchell, 2023) — inventing control structures or APIs that do not exist in the target programming language, borrowing features from other languages, often generating code that fails to type-check or even fails to parse.
This paper describes our initial investigations into combining LLMs and program verification. Our main aim is to evaluate how well LLMs can synthesize correct Dafny code, including the formal specifications and validation conditions so that the code passes the Dafny verifier: as part of that effort, we are also interested in how the additional logical rigor required by verification enhances or degrades generation performance. This paper makes the following contributions:
  • We conduct the first empirical study of LLMs synthesizing verifiable Dafny methods.
  • We demonstrate that a prompt designed under the principles of Chain of Thought (CoT) with semantically similar few shot examples explaining how to decompose a problem step by step, can synthesize verified and correct Dafny methods with meaningful specifications for 58% of problems in our test dataset.
  • We release the dataset MBPP-DFY-153, a collection of 153 programming problems with specifications, solutions and tests in/for Dafny, based on the MBPP (Mostly Basic Python Programming) dataset curated by Google (Austin et al., 2021).
  • We demonstrate that program verification techniques can help ensure code generated by LLMs is correct.
Our conclusions lead to a number of actionable insights. Most obviously, our first contribution means that programmers developing Dafny systems could already consider using LLMs — indeed, for simple cases, an LLM may be able to generate both the specificiations of a Dafny method, and a verifiably correct method implementation that meets those specifications. For researchers in LLMs or verification, our second contribution, the dataset of 153 formally verified Dafny examples, should enable further development of prompts and models to improve the performance of LLMs.
Ultimately, synthesis of verified methods (let alone whole programs) is a huge problem space consisting of multiple sub-problems: this paper is an initial exploration of this space. Whereas our first two contributions speak to the end to end task of synthesizing specifications and implementations from a narrative prompt, our third contribution highlights that LLMs may prove more effective at tackling sub-parts of the problem: generating just specifications from narrative prompts, or code from specifications (with or without other prompts), generating parts of method bodies, or generating only verification annotations, such as loop invariants or termination conditions.
Taken together, these results lead to two intriguing speculative ideas: that efforts on LLMs for code should concentrate on generating verifiable programs; and that our approach of combining LLMs with program verification techniques should transfer to other problem domains.
2. Background & Motivation
2.1. Dafny: A Verification-Aware Programming Language
Dafny (Microsoft, 2023a; Leino, 2020) is a verification-aware statically typed programming language which was first developed in Microsoft Research (MSR) (Microsoft, 2023b); it is currently being developed with the support of the Amazon Automated Reasoning research group (Amazon, 2023). Dafny includes various features of several programming paradigms such as imperative, functional, and object oriented programming. Its toolchain includes translators for generating executable code in different target languages such as C#, Java, JavaScript, Go and Python (Dafny, 2023). But the distinguishing feature of Dafny is that it supports code verification via design by contract. For that, it follows the framework of Floyd-Hoare-style  (Hoare, 1969) program verification with preconditions, postconditions, loop invariants and other high-level formal proof synthesis features such as pure functions, predicates, lemmas, and automated proof by inductions.
To develop a verified program, developers write Dafny code along with the specifications (pre-and post-conditions), and then add loop invariants and assertions that help the Dafny verifier prove the correctness of their code. While coding in VS Code, the Dafny static program verifier checks the functional correctness based on developers’ defined specifications and annotations. Correctness means the absence of any runtime errors with respect to the formal pre- and post-conditions, which means that the code does what the developers specify it to do. In order to confirm that the developer
1 : method FindSmallest(s: array<int>) returns (min: int)
2 :     requires s.Length > 0  // preconditions
3 :     // postconditions
4 :     ensures forall i :: 0 <= i < s.Length ==> min <= s[i]
5 :     ensures exists i :: 0 <= i < s.Length && min == s[i]
6 : {
7 :     min := s[0];
8 :     for i := 1 to s.Length
9 :         invariant 0 <= i <= s.Length  //loop-invariants
10:         invariant forall k :: 0 <= k < i ==> min <= s[k]
11:         invariant exists k :: 0 <= k < i && min == s[k]
12:     {
13:         if s[i] < min { min := s[i]; }
14:     }
15: }
16: //Main program, with a test case (checked statically!)
17: method Main() {
18:     var a1:= new int[] [1,2,3];
19:     var out1:= FindSmallest(a1);
20:     assert out1 == 1;
21: }
Figure 1. An example Dafny method; FindSmallest
specifications holds, the Dafny program verifier first transforms the code into an intermediate verification representation (IVR) known as Boogie IVR (Le Goues et al., 2011) language that expresses the verification conditions into predicate calculus (Leino, 2017). Next, an SMT solver tries to prove the verification conditions. In this phase, the Dafny verifier is backed by an advanced SMT solver known as the Z3 theorem prover (de Moura and Bjørner, 2008). The validity of these verification conditions implies the correctness of the code under consideration (Leino, 2017). However, sometimes the Z3 solver cannot automatically reach the required proof, even though such proof exists. In such cases, the developer intervention is required to give more context in the specifications by writing auxiliary verification conditions in the form of assertions, functions, predicates, and lemmas.
In recent years, Dafny has had major successes: Microsoft used Dafny to formally verify security libraries and kernels (Klein et al., 2009; Erbsen et al., 2020), distributed systems (Hawblitzel et al., 2015), and concurrent programs (Hawblitzel et al., 2014); Intel is developing its hardware encryption library using Dafny (Yang et al., 2023); ConsenSys successfully applied Dafny for their Ethereum Virtual Machine (EVM) verification (Cassez et al., 2023); Amazon implemented the Amazon Web Service (AWS) authorization and encryption logic in Dafny and deployed the Dafny-generated Java code into production (Cook, 2018; Chakarov et al., 2022).
2.2. Motivating Example
To illustrate how a Dafny program is statically verified, Figure 1 shows a simple procedure to find, and return, the smallest number in a given input array. The method FindSmallest is formally specified not just by the types of input and output parameters, but also with preconditions (with requires clause) and postconditions (with ensures clause). The postconditions, in particular, are the properties that we want to ensure are true for this method – i.e. what it means for this method to be “correct.” The Z3 SMT solver is usually not powerful enough to infer the postconditions from functional code alone; as such, developers must give additional hints, or verifiable conditions, along with the code in order to help the solver prove correctness of the postconditions.
In this particular case, the code contains a loop that modifies state, so we can provide loop invariants (with invariant clause) that must hold in every iteration of the loop; those invariants
are then checked by the SMT solver and, if they hold, that helps proving the postconditions of the method. The Main method represents the entry point: it invokes the method FindSmallest for some test inputs and checks (with assert clause) the outputs. Like other specification clauses, the assert statements are also statically checked. In this example the verifier tries to prove the assertion based on the postconditions of the FindSmallest method. All specification clauses such as requires, ensures, invariants, and assert are used statically – these clauses are not compiled into the executable and, as such, do not cause any runtime overhead.
Writing the functional code in Dafny is similar to writing code in other programming languages. However, researchers observed that writing the formal specifications and auxiliary verification code in Dafny is still difficult (Faria and Abreu, 2023). For instance, in Figure 1, the FindSmallest method requires 6 lines of formal specifications to verify a simple 5 line of algorithmic code. This example substantiates that the difficulty of writing formal specifications exponentially increases with the difficulty of the verifying program (Noble et al., 2022).
2.3. Large Language Models (LLMs)
Large Language Models (LLMs) are a class of deep neural networks pre-trained in publicly available data to perform a variety of downstream tasks in natural language processing (NLP). In the past year, LLMs have significantly impacted a wide range of application domains including code related tasks such as code generation, comprehension, summarization, and transformation.
The recent improvement of transformer based models (Radford et al., 2018) such as GPT (Generative Pre-trained Transformer) advanced the usages of LLMs in many coding related tasks compared to other pre-trained models (Chen et al., 2023; OpenAI, 2023a). For instance, CodeX  (Mark Chen et al., 2021) is a descendant of the GPT-3 based model trained on massive publicly available code repositories in GitHub, and is capable of performing coding related tasks in various programming languages. Codex powers GitHub Copilot (Copilot, 2023), an AI-pair programmer designed to assist software developers. OpenAI (OpenAI, 2023b) has now discontinued its Codex APIs, and suggested researchers adopt the successors of GPT-3 (GPT-3.5 and GPT-4). Recently Google released PaLM-2 (Rohan Anil et al., 2023) a new successor of their earlier PaLM (Aakanksha Chowdhery et al., 2022) model which is very capable of code generation and reasoning tasks.
Although many research studies have been published on the use of LLMs in many coding-related tasks, very few of them explore these models’ capability for synthesis of formally verified programs (First et al., 2023). One of the obstacles is that languages for formal code verification are not very popular. Hence, the training data available for LLMs to learn these languages, and their underlying concepts, pales when compared to the amount of training data available for popular languages such as Python, Java, C++, etc. Nevertheless, LLMs have been found to be capable of working as generic reasoning engines, including having the capability to solve complex problems and prove theorems (Li et al., 2023; Wu et al., 2022) that motivates us exploring how LLMs with the combination of theirs reasoning, theorem proving, and coding abilities can contribute to synthesizing verified Dafny code, along with formal specifications and validation conditions.
2.4. Prompt Engineering
Prompt engineering is a relatively new term and concept that emerged in the context of LLMs. It denotes the practice of iteratively crafting prompts for improving the generative power and fitness of purpose of an LLM. First coined by Gwern Branwen in 2020, [^3] https://gwern.net/gpt-3 prompt engineering has become a thriving subfield of research in LLMs (Bach et al., 2022; Dang et al., 2022; Deckers et al., 2023; Hou et al., 2022; Jiang et al., 2022b; Liu and Chilton, 2022).
LLMs are nominally intuitive to interact with: we give them a prompt, and they reply with a related text or image. The resulting output of the model relates to our prompt in plausible form. However, when using LLMs for specific technical goals, the simplistic approach of asking for something informally usually ends up in disappointing results.
Prompt engineering involves a systematic approach of creating prompts that effectively convey the desired downstream task to the pre-trained LLMs (Reynolds and McDonell, 2021; Beurer-Kellner et al., 2023). The crafting of prompts plays a significant role in the performance of LLMs retrieved contents. The prompt template should be structured in a manner that aligns with both the targeted task and the selected LLMs. Prompting techniques can be employed to train the LLMs explicitly through few-shot learning, incorporating a limited number of examples, such as method signatures and test cases. In the context of formally verified programs synthesis, the prompts need to be tailored for the targeted programming language (in our case Dafny) with unambiguous task descriptions (Bach et al., 2022; Dang et al., 2022).
3. Study Design
This study aims to investigate the current capability and potentiality of LLMs at synthesizing formally verified methods in Dafny. Specifically, we seek to answer the following research questions.
  • RQ1 [Contextless Prompting]: How effective are LLMs at synthesizing formally verified Dafny methods from simple natural language problem descriptions?
  • RQ2 [Signature Prompting]: How does additional context (signatures and tests) affect the synthesis outcome of LLMs?
  • RQ3 [Dynamic Few-Shot Prompting]: Do prompts designed with Chain of Thought (CoT) and semantically similar few-shot examples explaining problem decomposition step by step improve the synthesis outcome of LLMs?
3.1. Test Dataset: MBPP-san-DFY
In order to conduct our study, we needed a collection of problems that has natural language problem description corresponding formally verified Dafny code. We searched Dafny code in GitHub that yielded only 188 repositories having no natural language specifications of the available code. We would have to recover the informal natural language specifications from the formal code but by doing so we might inject errors by using words with a different intention than the original developers, or inject our own biases.
Since we aim to synthesize Dafny code from natural language specifications , for this purpose, we adopted the Mostly Basic Programming Problems (MBPP) benchmark published by Google Research in 2021 (Austin et al., 2021), which was used in prior program synthesis and verification studies. MBPP consists of a collection of close to 1,000 crowd-sourced algorithmic tasks with text descriptions, function signatures, concise implementations in Python, and three test cases to check their functional correctness. The tasks focus on basic arithmetic, array and integer sequences, and string processing. The textual descriptions of these tasks are concrete enough that a human should be able to translate them into code without any additional clarifications.
Solutions to 427 MBPP tasks have been inspected manually, and published as a subset called MBPP-sanitized (MBPP-san). We based our data on MBPP-san. However, since our goal is to synthesize Dafny, we chose to remove tasks that relied on Python-specific data types or external libraries, such as regex, math, heapq, etc.  [^4] Dafny does not have a standard library, much less external libraries. Note that any Dafny libraries would need to be formally-verified a prior, or else the Dafny tasks would become much more complex than their Python versions. We culled 82 tasks, leaving a dataset of 345 problems.
In order to keep our study’s costs under budget, we randomly selected 228 problems. Out of these 228 problems, we picked a random sample of 50 to act as verified Dafny exemplars (see below Section 3.2) leaving us with 178 remaining problems for testing. We manually translated each of those 178 problem descriptions, signatures, and test cases into Dafny, but we did not translate the solutions’ Python code, or attempt to write specifications. This resulted in our main test dataset, MBPP-san-DFY: 178 instances of MBPP-san problems, each one with a description, method signature, and three test cases translated to suit Dafny.
The MBPP problem set, and Python solutions to the MBPP problems, are publicly available and it seems likely they would have been incorporated into the LLMs’ training sets. However, there are no Dafny implementations of MBPP. Therefore, we hypothesize that employing MBPP-san-DFY dataset would not create any data leakage in our experiment.
3.2. Human Written Dataset: MBPP-DFY-50
To perform Dynamic Few-Shot prompting, we needed a relatively large and varied collection of verified Dafny methods that could serve as examplars to be included in prompts. A search for an existing suitable collection proved fruitless. Our best alternative was to translate a subset of MBPP-san into Dafny ourselves, but now also translating the Python solution for each problems into Dafny, then coming up with specifications (method pre- and post-conditions, loop invariants) for each example, and finally doing whatever annotation, correction, or rewriting was needed until the Dafny verifier would signal that the method was correct.
In this process, we experienced first-hand how hard it is to formally specify postconditions and to provide loop invariants and hints so that correct code can pass the Dafny verifier. It took approximately 220 hours for the first two authors to create this dataset of 50 formally verified problems, even with the Python solutions, full access to Dafny documentation, books, blog posts by experts, and Stack Overflow.
3.3. LLM Selection
Program synthesis in Dafny requires knowledge of Dafny syntax, as well as reasoning and theorem-proving abilities (Noble et al., 2022; First et al., 2023). Therefore, we looked for LLMs that possess a combination of these abilities. To determine which LLMs to use, we performed a pilot study using our handwritten dataset MBPP-DFY-50. We started by looking at GPT-4 (OpenAI, 2023a) and GPT-3.5 (Mark Chen et al., 2021) from Open-AI (OpenAI, 2023b), PaLM-2 (Rohan Anil et al., 2023) and PaLM (Aakanksha Chowdhery et al., 2022) from Google, and CodeT5+ (Wang et al., 2021), CodeT5 (Wang et al., 2023) from Salesforce and Code-Llama (Rozière et al., 2023) from Meta. All these LLMs have been explored in prior code generation, program synthesis, and reasoning-related studies. Our pilot study consisted of querying all these LLMs with the problem description, asking them to solve each. We then analyzed their responses to get a sense of their reasoning abilities, and the quality of their approaches compared to our ground truth. For these 50 problems, CodeT5, CodeT5+, LLaMA, Code-LLaMA, and PaLM were not able to synthesize a single syntactically correct Dafny method with specifications, as all of them encountered compilation errors. Even in GPT-3.5, we got only 5/50 syntactically correct solutions. However, GPT-4 and PaLM-2 generate verified Dafny methods in 16/50 and 7/50 problems respectively. Based on this, we decided to narrow our investigation to GPT-4 and PaLM-2 only.
3.4. Prompt Design
Based on the background literature and our pilot study, we knew that our experiment would depend on the prompts we offered to the LLMs. Therefore, we experimented with three different prompts.
3.4.1. RQ1 [Contextless]
To answer RQ1 we designed a simple prompt that identifies the programming language (Dafny), the problem description, and instructions for how to format the code response. Our prompt design is inspired by Austin et al. (Austin et al., 2021). This kind of prompt works well for popular languages like Python, JavaScript, Java, C#, and C++.
You are an expert Dafny programmer, and here is your task: Write a method in Dafny to find maximum run of uppercase characters in the given string. You must return the method in the following form: ```dafny //Dafny Code ```
You are an expert Dafny programmer, and here is your task: Write a method in Dafny to find maximum run of uppercase characters in the given string. Method Signature: method maxRunUppercase (str: string ) returns (res: int ) Your code should satisfy these tests: var out1:=maxRunUppercase( "GeMKSForGERksISBESt" ); assert out1==5; var out2:=maxRunUppercase( "PrECIOusMOVemENTSYT" ); assert out2==6; var out3:=maxRunUppercase( "GooGLEFluTTER" ); assert out3==4; You must return the method in the following form: ```dafny //Dafny Code ```
Figure 2. An example of Contextless and Signature prompt
The Contextless prompt instructs the LLMs to synthesize a Dafny method in a single step without providing any additional context. Figure 2 (top) shows an example of Contextless prompt. Since writing functional code in Dafny is similar to writing code in high-level programming languages, with Contextless prompt our goal is to establish a baseline of the LLMs’ knowledge of Dafny.
3.4.2. RQ2 [Signature]
For RQ2, we extended the Contextless prompt with the desired method’s signature and three test cases (see Figure 2 (bottom)). Code generation experiments for non-verified code often include test cases and signatures (implicitly or explicitly) (Austin et al., 2021; Mark Chen et al., 2021). We hypothesized that with the method signature and test cases LLMs can infer the pre- and post-conditions that would be beneficial to synthesize complex problems.
3.4.3. RQ3 [Dynamic Few-Shot]
Work in LLMs has shown that Chain of Thought (CoT) reasoning to decompose a problem into multi-intermediate steps can lead to drastically improved reasoning, especially for code synthesis tasks (Madaan et al., 2022; Wei et al., 2022; Reynolds and McDonell, 2021; Beurer-Kellner et al., 2023).
Used together with relevant examples in the prompt, aka few-shot prompting, it becomes possible to “teach” LLMs things that had a weak presence in the training data. Inspired by this, we designed a prompt following the principle of Chain of Though (CoT) with semantically similar few-shot Dafny method examples where each example explains its problem decomposition step by step and how to derive the implementation in Dafny. The thought process and reasoning steps of this prompt was informed by our own experience with Dafny. When we write Dafny methods, we start by defining the method’s signature, followed by thinking about the pre- and post-conditions that should hold for that signature. Then we write the functional code, i.e. the algorithm that we believe should result in those postconditions to hold, and finally we add the proof, i.e. the necessary loop invariants and verification annotations that lead the Dafny verifier to conclude that the algorithm is, indeed, correct.
However, for a given problem, this prompt will be effective when semantically similar examples of the given problem are included in this prompt. Therefore, we sought for a heuristic that can be employed to search best examples for a given problem. While creating our MBPP-DFY-50 dataset, we observed that the postconditions are the most valuable parts of Dafny method specifications. For each problem we tackled, we searched through the problems we had already solved, looking for similar postconditions. We then used those examples to guide the proof of the new problem. We identified a pattern that problems having similar method signature and post-conditions require the same type of Dafny proofs and code to solve the problem. This pattern motivated us to develop a prompt chaining pipeline for selecting few-shot Dafny code examples that have similar method signatures and post-conditions for a given problem. Figure 3 depicts the high-level overview of our developed pipeline for prompt chaining.
Figure 3. A high-level overview of the prompt-chaining pipeline for Dynamic Few-Shot Prompting.
To select semantically similar Dafny methods based on postconditions, we employed prompt chaining in two phases. In the first phase, for a given problem, we simply prompt the LLM to generate the method specifications such as signature and postconditions. To teach LLM how to generate such specifications, we provided three few-shots examples of problem description, method signature and postcondition pairs from our MBPP-DFY-50 dataset. We elected these examples based on the semantic similarity of the given problem description. An example of such specification prompting is shown in Figure 4 (top).
TASK: Write a method in Dafny to convert a given string to uppercase. Method Signature: method ToUppercase(s: string ) returns (v: string ) Postconditions: - The length of the returning string must be the same as the length of the input string - All lowercase characters of the input string are uppercase in the output string - All other characters are unchanged ------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------ //Two more semantically similar few-shots example TASKS ------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------ TASK: Write a method in Dafny to find maximum run of uppercase characters in the given string.
SYSTEM: You are an expert AI assistant that writes Dafny programs. You are very good at writing verifiable correct code in terms of preconditions and postconditions of methods, and at finding the appropriate loop invariants for the pre/postconditions to hold. ------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------ TASK: Write a method in Dafny that matches a word containing ’z’. AI ASSISTANT: Step 1: Determine the method’s signature. In this case: method ContainsZ(s: string) returns (result: bool ) Step 2: Determine the preconditions. In this case: - There are no preconditions, the method will work. Strings and sequences are always not null. Step 3: Determine the postconditions. In this case: - The result is true if the input string contains the character ’z’ or ’Z’ - The result is false if the input string does not contain the character ’z’ or ’Z’ Step 4: Write the code, its formal specification, and the verification annotations that help the solver prove the postconditions. In this case: ‘‘‘dafny     method ContainsZ(s: string) returns (result: bool)         ensures result <==> (exists i :: 0 <= i < |s| && (s[i] == ’z’ || s[i] == ’Z’))     {         result := false;         for i := 0 to |s|             invariant 0 <= i <= |s|             invariant result <==> (exists k :: 0 <= k < i && (s[k] == ’z’ || s[k] == ’Z’))         {             if s[i] == ’z’ || s[i] == ’Z’ {                 result := true;                 break;             }         }     } ‘‘‘ ------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------ //Four more semantically similar few-shot example TASKS ------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------ TASK: Write a method in Dafny to find maximum run of uppercase characters in the given string. AI ASSISTANT:
Figure 4. An example of Specification Prompt (top) and CoT Prompt (bottom)
In the next phase, we search for semantically similar Dafny methods in our MBPP-DFY-50 dataset, incorporating the LLM generated specification response. To perform the semantic similarity, we utilized OpenAI Embedding API that computes semantic similarity based on cosine distance. For the given specification generated in the first phase, we selected top five Dafny methods examples from our MBPP-DFY-50 dataset. We used 5 examples so that we can fit the whole prompt in LLMs’ context length (i.g., token limits). With these few-shot examples, we then prepared a CoT prompt that first instruct LLMs to synthesize the specifications according the problem decomposition steps and then synthesize a Dafny method that satisfies these specifications. Figure 4 (bottom) depicts an example of our designed CoT prompt.
3.5. Evaluation Metrics
We evaluate LLMs’ synthesized verified methods against the Dafny verifier. The verifier reports whether Dafny code is (a) syntactically correct and (b) verified. We report the synthesis performance using the pass@k metric (Kulal et al., 2019; Mark Chen et al., 2021; OpenAI, 2023a; Wang et al., 2023). In our study, we use verify@k metric (similar to pass@k), where k sample methods are synthesized per problem; a problem is considered to be verified if any attempt verifies.
Automated quantitative metrics are only a weak proxy for performance of LLMs in this task. In one extreme, the synthesized Dafny code may not contain any formal specifications whatsoever; in those cases, the verifier will report verification success. Synthesized methods may contain only weak and uninteresting formal specifications that are trivial to verify. The most challenging part of this study is to check whether LLMs are able to generate the “right” and “strong” formal specifications — in particular, postconditions. In order to assess the semantic correctness of the generated verification code, we manually reviewed all verified methods. We also manually inspected all the generated methods that failed to verify, to identify what types of errors prevented verification.
3.6. Temperature Tuning
The “temperature” is an important parameter of all LLMs that controls the randomness of the decoding process (OpenAI, 2023a; Rohan Anil et al., 2023). In doing experiments with LLMs, it is important to identify the best temperature for each experiment. Therefore, before running our experiments in the entire MBPP-san-DFY dataset, we first performed a temperature tuning experiment. We chose a random sample of 50 problems from our test dataset MBPP-san-DFY of 178 problems employing the sampling procedures with 90% confidence interval and 10% margin of error. Next, for each type of prompt, we executed these 50 problems in four different temperatures $T\in\{0.0,0.25,0.5,0.75\}$ in both GPT-4 and PaLM-2 . We evaluated the synthesized Dafny methods with our evaluation metrics verify@k where $k\in\{1,3,5\}$. As shown in Figure 5 , we observed that in Contextless and Signature prompt, GPT-4 shows best performance in T=0.75. However, in Dynamic Few-Shot prompt, GPT-4 is most effective in T=0.5. PaLM-2 demonstrates better performance with T=0.0 in Contextless prompt. And in Signature and Dynamic Few-Shot Prompts PaLM-2 works good in T=0.5. Thus, we used those values for GPT-4 and PaLM-2 , and their respective prompts.
Figure 5. Temperature tuning for GPT-4 and PaLM-2 on a sample of 50 problems in four different temperatures $T\in\{0.0,0.25,0.5,0.75\}$ for three different prompts. The values in vertical axis represents the verified program synthesis rate in verify@k (k=5).
3.7. Experimental Setup
We developed scripts to execute different prompts on different temperatures through Open-AI’s API for GPT-4 and Google’s API for PaLM-2. For verifying the code, we employed the latest Dafny 4.0.0 version  [^5] https://dafny.org/blog/2023/03/03/dafny-4-released/. In March 3, 2023, a new major version, Dafny 4.0.0, has been released that broke backwards compatibility. Both GPT-4 and PaLM-2 may not know anything about this latest version. In our prompts and few-shot examples we did not specify which version of Dafny should be used. We verified our MBPP-DFY-50 dataset in Dafny 4.0.0. For executing the experiments and verifying synthesized Dafny methods, we used a MacBook Pro (13-inch, 2020), having 2 GHz Quad-Core Intel® Core™ i5 processor, 16GB RAM, 512 GB SSD and running macOS Ventura 13.4.
4. Results and Analysis
To evaluate LLM synthesis of Dafny methods, we queried all 178 problems in three prompts with their tuned temperatures (described in Section 3.6). However, we noticed that LLMs cannot generate any Danfy method snippets if they misunderstood the problem or had insufficient knowledge to solve those problems. In such cases, LLMs returned a textual explanation why it cannot generate any Dafny method snippets: we filtered out those responses. We then analyzed the verification results quantitatively and qualitatively.
4.1. Quantitative Analysis
In this section, we explain and compare our quantitative analysis across three different prompts for GPT-4 and PaLM-2 .
4.1.1. Verification:
Table 1 summarizes the overall verification results for three prompts queried with their tuned temperatures. For example, in Contextless prompting at k=5, GPT-4  synthesized 104 verified methods out of 178 problems, achieving a verification success rate 58.42%. We notice that the verification success rate improves with the increasing number of attempts. For instance, in Contextless for GPT-4 in k=1, the verification success rate is 32.58% whereas at k=3 and k=5 it jumps to 53.37% and 58.42% respectively. PaLM-2 was unable to generate a single verified method for Contextless prompting even with several attempts. In Signature prompting, the verification success rate also improves with the number of attempts. Here, GPT-4  generated 95 Dafny verified methods (53.37%), which is slightly lower than Contextless prompting. PaLM-2   gets a boost at k=5, synthesizing 12 (6.74%) verified methods. Both GPT-4 and PaLM-2 attain the best performance in Dynamic Few-shot prompting. GPT-4 got 114 verified methods (64.04%) in k=5 that surpass Contextless prompting’s performance. PaLM-2 achieved significant improvement with a 29.21% verification success which is around 4X performance boost compared to Signature prompting.
Table 1. Summary of verified Dafny method synthesis on 178 problems. For each prompt, LLMs responses were generated with their tuned temperature (T) in verify@k (where $k\in\{1,3,5\}$)
Prompts GPT-4  PaLM-2
Contextless Temperature verify@k Temperature verify@k
T=0.75 k=1 58/178 (32.58%) T=0.00 k=1 0/178 (0.00%)
k=3 92/178 (51.64%) k=3 0/178 (0.00%)
k=5 104/178 (58.42%) k=5 0/178 (0.00%)
Signature Temperature verify@k Temperature verify@k
T=0.75 k=1 59/178 (33.14%) T=0.50 k=1 12/178 (6.74%)
k=3 88/178 (49.43%) k=2 9/178 (5.05%)
k=5 95/178 (53.37%) k=3 12/178 (6.74%)
Dynamic Few-Shot Temperature verify@k Temperature verify@k
T=0.50 k=1 86/178 (48.31%) T=0.50 k=1 30/178 (16.85%)
k=3 105/178 (58.98%) k=3 43/178 (24.15%)
k=5 114/178 (64.04%) k=5 52/178 (29.21%)
4.1.2. Postconditions:
In a naive interpretation, these results could indicate that the complicated third prompt is not worth it, at least for GPT-4 . However, that is far from being the case. Verification success does not mean that the solution was properly formulated. Postconditions, specifically, are essential in formally specifying what methods are supposed to do. The absence of postconditions is a strong sign that the solution is not usefully verifiable: methods with no specifications trivially avoid being wrong.
Table 2. Summary of specifications synthesis result, in both verified and unverified methods. For each prompt, LLMs responses were generated with their tuned temperatures at verify@k (k=5).
Prompts GPT-4  PaLM-2
Contextless Specifications Verified Unverified Verified Unverified
has postconditions 56/104 (53.84%) 48/73 (65.75%) N/A 104/176 (59.09%)
has preconditions 64/104 (61.53%) 47/73 (64.38%) N/A 2/176 (1.13%)
has both 41/104 (39.42%) 32/73 (43.83%) N/A 2/176 (1.13%)
has invariants 40/104 (38.46%) 28/73 (38.35%) N/A 4/176 (2.27%)
Signature Specifications Verified Unverified Verified Unverified
has postconditions 30/95 (31.57%) 22/80 (27.5%) 1/12 (8.33%) 62/162 (38.27%)
has preconditions 46/95 (48.42%) 31/80 (38.75%) 1/12 (8.33%) 61/162 (37.65%)
has both 21/95 (22.10%) 19/80 (23.75%) 1/12 (8.33%) 61/162 (37.65%)
has invariants 38/95 (40.00%) 19/80 (23.75%) 0/12 (0.00%) 6/162 (3.70%)
Dynamic Few-Shot Specifications Verified Unverified Verified Unverified
has postconditions 114/114 (100%) 61/61 (100%) 52/52 (100%) 123/123 (100%)
has pre conditions 74/114 (64.91%) 42/61 (68.85%) 35/52 (67.30%) 99/123 (80.48%)
has both 74/114 (64.91%) 42/61 (68.85%) 35/52 (67.30%) 99/123 (80.48%)
has invariants 57/114 (50.0%) 57/61 (93.44%) 13/52 (25.0%) 90/123 (73.17%)
For both the verified and unverified methods, we looked into the formal specification synthesis results. In Contextless prompting as shown in Table 2, we noticed that, although, GPT-4  generated 104 verified methods, only 56 of them contain postconditions. Since the remaining 48 verified methods have no postconditions, indeed, they cannot be considered formally verified. Regarding unverified methods, GPT-4 generates postconditions in 65.75% of them. Similar to GPT-4 , PaLM-2 also shows better effectiveness (59.09%) in synthesizing postconditions in unverified methods. However, since these methods are unverified, it may happen that either these postconditions are wrong or the syntactic or semantic implementation of these methods is incorrect to prove the correct postconditions. In our qualitative analyses we manually examine these methods and their specifications to identify the actual reasons (see Section 4.2).
In Signature prompting, the presence of postconditions gets worse. Only 30 out of 95 synthesis methods from GPT-4 have postconditions whereas for PaLM-2 only 1 method out of 12 contains postconditions. For unverified methods the presence of postconditions drastically drops for both GPT-4 (31.57%) and PaLM-2 (38.27%). It indicates that including additional contexts in the prompt did not bring any substantial benefit.
However, in Dynamic Few-Shot prompting, the specification synthesis results outstandingly improves for both GPT-4 and PaLM-2 . According to Table 2, 100% of both verified and unverified methods have postconditions. The primary reason for having postconditions in all synthesized methods is the few-shot examples with the CoT process. We designed this prompt to dynamically select five few-shot example methods based on their postconditions’ semantic similarity, and each of these examples includes a CoT process describing how to determine the post-conditions in natural instructions and then in Dafny method in the implementation. The design of this prompt made the LLMs follow the manner by which the examples were implemented.
4.1.3. Preconditions:
In a Dafny method, preconditions are only required when certain properties or assumptions must be hold before invoking the method. Having preconditions is not a mandatory formal specification that needs to be proved by the methods’ implementation.
In our experiment with Contextless prompting, we found that 61.53% of GPT-4 generated verified methods contain preconditions and 39.42% contain both pre-and postconditions (Table 2). Like postconditinos, the presence of preconditions drops in Signature prompting where only 48.42% of verified methods contain preconditions and 22.10% have both pre-and postconditions. Unfortunately, in Contextless and Signature prompting, PaLM-2 specification synthesis performance is trivial. However, for both GPT-4 and PaLM-2 , Dynamic Few-Shot prompting enhances the specification synthesis performance. As demonstrated in Table 2, on average 64% of verified and unverified programs for both GPT-4 and PaLM-2 contain pre-and postconditions. It substantiate that the design of Dynamic Few-Shot prompt considerably strengthens LLMs specification synthesis performance.
4.1.4. Invariants:
When a Dafny method has a loop, it should contain invariants that need to be maintained in each iteration of the loop. As shown in Table 2, in Contextless prompt, on an average 38% GPT-4 synthesis methods contain invariants. In Signature prompting, the number increases slightly for verified methods, however, the presence of invariants exceptionally enhances for Dynamic Few-Shot prompting in GPT-4 synthesis methods. According to Table 2, 50% of verified and 93.44% of unverified methods contain at least one loop invariant. On the contrary, PaLM-2 struggles to synthesize loop invariants for both verified and unverified methods in Contextless and Signature prompting. Only the Dynamic Few-Shot prompting outstandingly improves its performance to have loop invariants in 25% of verified and 73.17% for unverified methods respectively.
4.2. Qualitative Analysis
This section describes our manual investigation. The first two authors manually examined all prompt responses, assessing the quality of the formal specifications, specifically the postconditions of the verified Dafny methods – again, postconditions are the key element for a correct formal specification. During this manual inspection, when we have any doubt about the implementation or specifications, we looked into our ground-truth dataset, MBPP-san, and discussed with each other. The two authors independently performed this manual tagging and reached a high Inter-Rater Reliability (IRR) (Hallgren, 2012)(93.91% Cohen Kappa (Viera et al., 2005)), which ended up with consensus at the end. We summarize our qualitative analysis findings in Table 3. The complete analysis results are available in our companion repository 8, including all prompts, LLMs’ responses, verification results, Dafny method snippets with our manual tagging.
Table 3. Summary of qualitative analysis of specifications synthesis in verified methods. For each prompt, LLMs responses were generated with their tuned temperatures at verify@k (k=5).
Specifications Groups GPT-4 PaLM-2
Prompts Prompts
Contextless Signature
Dynamic
Few-Shot
Contextless Signature
Dynamic
Few-Shot
Postconditios Strong 34/56 18/30 103/114 (90.35%) N/A 1/1 35/52
Weak 15/56 9/30 3/114 N/A 0/1 3/52
Wrong 7/56 3/30 8/114 N/A 0/1 14/52
Preconditions Required 19/41 10/21 36/74 (48.64%) N/A 0/1 6/35
Optional 22/41 11/21 38/74 N/A 1/1 29/35
Invariants Strong 10/27 5/16 51/57 (89.47%) N/A N/A 13/13
Weak/Wrong 17/27 11/16 6/57 N/A N/A 0/13
4.2.1. Postconditions:
The goal of a postcondition is to make sure that the implementation is correct based on the proofs. Here the term ”correct” implies that the postcondition is strong enough to precisely capture all possible output values. We tagged the quality of postconditions into three groups, Strong, Weak and Wrong. In Signature prompting, PaLM-2 got only one verified method and that has Strong postconditions. With regard to, Dynamic Few-Shot prompt, we found that in 14 cases, PaLM-2 misunderstood problems’ description and generated wrong method implementation for those problems. However, it got Strong postconditions in 35 methods and Weak postconditions in 3 methods only. Although GPT-4 got 56 verified programs in Contextless prompting, we encountered 15 of them have Weak postconditions. Table 4.2.1 exhibits such an example of Weak postconditions. On the other hand in Signature prompting 9 of verified methods contain Weak postconditions which are very trivial specifications to verify correct implementation.
In Dynamic Few-Shot, GPT-4 ’s specification synthesis result is impressive. Only 3 of verified methods got Weak postconditions and in 8 cases it misunderstood the problems’ description generated wrong methods. One interesting example of such cases where GPT-4 misunderstood the problem description and generated Wrong method is illustrated in Table 4.2.1. Specifically in Dynamic Few-Shot, GPT-4 got 90.35%(103/114) of verified methods containing Strong postconditions.
Table 4. Examples of LLMs synthesis Strong, Weak and Wrong postconditions; Required and Optional preconditions; invariants subsuming postconditions, and modularization of proofs.
Specifications LLM Synthesis Specifications in Dafny
Post- conditions Strong
Weak
/*Write a method in Dafny to interleave 3 sequence of the same
length into a single flat sequence.*/
method SplitAndAppend(l: seq<int>, n: int) returns (r: seq<int>)
    requires n >= 0 && n < |l|
    ensures |r| == |l|
    ensures forall i :: 0 <= i < |l| ==> r[i] == l[(i + n) % |l|]
Wrong
//Write a method in Dafny to find the median of two sorted lists of same size.
method FindMedian(a: array<int>, b: array<int>) returns (median: int)
    //...MISUNDERSTANDING
    ensures median == if (a.Length % 2 == 0)
                        then (a[a.Length / 2 - 1] + b[0]) / 2 else a[a.Length / 2]
Pre- conditions Required
//Write a method in Dafny to compute the n-th power of each number in a list.
function Power(base: int, exponent: int): int //.. PROOF MODULARIZATION
    requires exponent >= 0
    { if exponent == 0 then 1 else base * Power(base, exponent-1) }
method PowerOfListElements(l: seq<int>, n: int) returns (result: seq<int>)
    requires n >= 0
    ensures |result| == |l|
    ensures forall i :: 0 <= i < |l| ==> result[i] == Power(l[i], n)
Optional
//Write a method in Dafny to append the given array to the given sequence.
method AppendArrayToSeq(s: seq<int>, a: array<int>) returns (r: seq<int>)
    requires a != null
    /*In Dafny arrays and sequences are never null,
    unless the method declares them as array? and seq?*/
Invariants Subsume Post- conditions
/* Write a method in Dafny to check whether a list of sequence contains
the given sequence or not.*/
method ContainsSequence(list: seq<seq<int>>, sub: seq<int>) returns (result: bool)
    ensures result <==> (exists i :: 0 <= i < |list| && sub == list[i])
    result := false;
    for i := 0 to |list|
        invariant 0 <= i <= |list|
        invariant result <==> (exists k :: 0 <= k < i && sub == list[k])